OpenStudy (dls):

Derivative of (x^3+y^3)xy=(x^5-y^3)

OpenStudy (dls):

\[\large (x^3+y^3)xy=(x^5-y^3) \]

OpenStudy (dls):

Solve for dy/dx

OpenStudy (dls):

@zepdrix

OpenStudy (dls):

\[\large (x^4y+xy^4)=(x^5-y^3) \] like u said^^

OpenStudy (dls):

\[\large 4x^3y+x^4 \frac{dy}{dx}= x^4-3y^2 \frac{dy}{dx}\]

OpenStudy (dls):

\[\frac{dy}{dx}(x^4+3y^2)=x^4-4x^3y\]

OpenStudy (dls):

\[\large \frac{dy}{dx}=\frac{x^4-4x^3y}{x^4+3y^2} =\frac{x^3(x-4y)}{x^4+3y^2}\]

OpenStudy (dls):

sup sensei :D

zepdrix (zepdrix):

\[\large \left(4x^3y+x^4 \frac{dy}{dx}\right) + \qquad \qquad \qquad = 5x^4-3y^2 \frac{dy}{dx}\] Hmm it looks like you applied the product rule to the first term, what about the other one? :O

OpenStudy (dls):

ive applied it,check the next steps!

OpenStudy (dls):

@zepdrix ?? :(

zepdrix (zepdrix):

I dunno, I can't follow your steps :\

OpenStudy (dls):

okay..lets move!